OpenStudy (anonymous):

math (algebra 1) ASAP

zepdrix (zepdrix):

When we divide terms of similar base, we `subtract` the exponents. \[\large \frac{7^4}{7^2} \qquad = \qquad 7^{4-2}\] Can you solve it from there? :)

OpenStudy (anonymous):

I think it is =, because 7^4 - 7^2 = 7^2. 7^2 = 7 x 7, which equals 49, so i think the answer is = to.

zepdrix (zepdrix):

Yay good job :)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

7^4 = 2401 7^2 = 49 2401/49 = 49 49 = 49

OpenStudy (anonymous):

thanks!!!!!